OpenStudy (priyar):

If an equation of a tangent to the curve,y=cos(x+y), -1<= x <= 1+ π is x+2y=k then k is equal to?

OpenStudy (priyar):

i differentiated the curve eq to get the tangent eq.

OpenStudy (priyar):

then i equated the slope to -1/2 (i found this thro' the tang. eq)

OpenStudy (baru):

then find at what value of (x,y) you get slope= -1/2

OpenStudy (priyar):

i got x+y=π/2

OpenStudy (priyar):

after that how to find the value of x?

OpenStudy (priyar):

we need another eq for that...

OpenStudy (priyar):

should i substitute x=1+π or x=1?

OpenStudy (priyar):

*-1

Parth (parthkohli):

\[y' = -\sin(x+y) (1+y')\]But you're given \(y' = -1/2\). So \(x+y = \pi/2\). Therefore, \(y = \cos(\pi/2) = 0\). We need the point where \(y = 0\) and \(x + y = \pi/2\) intersect. So the point is \((\pi/2, 0\). Now substitute to the equation of tangent and you get \(k = \pi/2\). I'm not sure if we have other solutions to this.

OpenStudy (priyar):

why y=cos(π/2)=?

Parth (parthkohli):

y = cos(x+y) is your function

OpenStudy (priyar):

oh ya fine!

OpenStudy (priyar):

thank u